OpenStudy (anonymous):

A stone is dropped from the top of a cliff. It is seen to hit the ground below after 4.2 s. How high is the cliff?

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

we must first find the velocity of the stone after 4.2 seconds v=u+gt v=0+9.8(4.2) final velocity of the object after 4.2 seconds is 41.16m/s v^2=u^2+2gS s=v^2-u^2/2g s=(41.16)^2-(0)^2/2(9.8) s=86.43 m the height of the cliff is 86.44 metres above the ground